OpenStudy (anonymous):

LOGARITHMS. 12^(x-2)=20 Round To The Nearest Ten-Thousandths

OpenStudy (aum):

\(\large 12^{x-2} = 20\) Take logarithm on both sides: \(\large \log(12^{x-2}) = \log(20)\) \(\large (x-2)\log(12) = \log(20)\) Can you solve for x?

OpenStudy (anonymous):

Doesnt It Become Like 2xlog12=log20?